# EgoEvGesture: Gesture Recognition Based on Egocentric Event Camera

### v1

@@
EgoEvGesture: Gesture Recognition Based on Egocentric Event Camera
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
Egocentric gesture recognition is a pivotal technology for enhancing natural human-computer interaction, yet traditional RGB-based solutions suffer from motion blur and illumination variations in dynamic scenarios. While event cameras show distinct advantages in handling high dynamic range with ultra-low power consumption, existing RGB-based architectures face inherent limitations in processing asynchronous event streams due to their synchronous frame-based nature. Moreover, from an egocentric perspective, event cameras record data that include events generated by both head movements and hand gestures, thereby increasing the complexity of gesture recognition. To address this, we propose a novel network architecture specifically designed for event data processing, incorporating (1) a lightweight CNN with asymmetric depthwise convolutions to reduce parameters while preserving spatiotemporal features, (2) a plug-and-play state-space model as context block that decouples head movement noise from gesture dynamics, and (3) a parameter-free Bins-Temporal Shift Module (BSTM) that shifts features along bins and temporal dimensions to fuse sparse events efficiently. We further build the EgoEvGesture dataset, the first large-scale dataset for egocentric gesture recognition using event cameras. Experimental results demonstrate that our method achieves 62.7% accuracy in heterogeneous testing with only 7M parameters, 3.1% higher than state-of-the-art approaches. Notable misclassifications in freestyle motions stem from high inter-personal variability and unseen test patterns differing from training data. Moreover, our approach achieved a remarkable accuracy of 96.97% on DVS128 Gesture [1], demonstrating strong cross-dataset generalization capability. The dataset and models are made publicly available at https://github.com/3190105222/EgoEv_Gesture.
1 Introduction
@@
To address these challenges, we first propose a novel approach with three progressive processing stages, as illustrated in Figure 6. First, we convert raw events into Locally Normalized Event Surfaces (LNES) [13] that focus on the most recent temporal signature, suppressing historical ego-motion artifacts while maintaining real-time performance. Second, we employ Blaze [14] & Vision Mamba [15] (VMamba) blocks to extract spatiotemporal features. The Blaze encoder-decoder, adapted from [14] with proven efficiency in human-centric perception [16], processes each temporal bin efficiently. The VMamba block  [15] addresses sparse event processing and long-term temporal modeling through discrete state space equations. Finally, we propose the Bins-Temporal Shift Module (BTSM), a zero-parameter mechanism that shifts features bidirectionally along two orthogonal axes: 1) Bins axis (fine temporal resolution: 33 ms/bin) for inter-bin context fusion, 2) Raw temporal axis (coarse resolution: 200 ms) for intra-bin motion alignment. This dual-scale shifting resolves temporal fragmentation while maintaining motion invariance. Due to the lack of event datasets in the egocentric setting, we developed EgoEvGesture, the first large-scale dataset for egocentric gesture recognition using event cameras. Captured under both normal and low-light conditions, it includes data from 10 subjects performing a variety of gestures, offering rich temporal and spatial information for training and evaluating models in real-world scenarios. Our proposed method achieves a classification accuracy of 62.7% on this dataset, with only 7M parameters, outperforming existing (closely related) approaches by 3.0%. To verify the effectiveness of our approach, we evaluated it on the DVS128 Gesture dataset [1], achieving a competitive accuracy of 96.97%. This result highlights the robustness of our architecture in handling sparse and temporally challenging event-based data. The significance of each proposed module is evaluated and confirmed in an ablation study.
In summary, this paper defines a new problem, i.e. gesture recognition based on egocentric event camera, and makes the following technical contributions:
@@
2 Related Work
Next, we review related works on egocentric gesture recognition and event-based approaches to human perception. As shown in Table 1, our EgoEvGesture represents the first dataset for egocentric gesture recognition that uses real-world event streams, offering several advantages over existing RGB- or depth-based methods, as highlighted in the following sections.
2.1 Egocentric Gesture Recognition
Egocentric gesture recognition using RGB and depth sensors has been extensively explored [17, 18, 19, 6, 20, 21, 22]. The EgoGesture dataset [7], one of the largest benchmarks, contains 83 gesture classes designed for wearable device interactions, which inspired our motion design. The EPIC-KITCHENS dataset [23] focuses on action classification in kitchen environments, but prioritizes general activities over specific gestures. The FPHA dataset [24] pioneered 3D hand pose annotations in egocentric scenarios but suffers from magnetic sensor interference and lacks two-hand pose data. The H2O dataset [25] is the first dataset to emphasize bimanual object interactions and achieves high precision in gesture recognition. However, it primarily relies on static image annotations, and lacks continuous dynamic gesture capture, which limits its applicability to real-time gesture recognition. Assembly101 [26] and ARCTIC [27] extend the scope of egocentric hand-object interaction datasets. While Assembly101 provides multi-view videos for activity understanding, its focus on assembly tasks restricts gesture diversity. ARCTIC offers dexterous bimanual manipulation data but lacks dynamic gesture annotations. HOI4D [28] introduces category-level human-object interaction labels, including panoptic segmentation, motion segmentation, and reconstructed 3D meshes, yet it only addresses single-hand manipulation tasks.
Traditional gesture recognition methods, such as SVM [29, 30], Random Forests [31], KNN [32, 33], and HMM [34, 35], rely on handcrafted features (e.g., optical flow, edge descriptors) to model gesture dynamics. These approaches are susceptible to noise in complex environments and struggle to capture subtle motion variations, resulting in degraded performance for fast or intricate gestures. Deep learning has become the mainstream for gesture recognition. 2D CNNs [36, 37] extract spatial features from single frames, whereas 3D CNNs [38, 39] capture short-term spatiotemporal dependencies. However, single CNN models often fail to balance global and local feature interactions in egocentric scenarios. RNNs [40, 41] model temporal relationships but face challenges such as vanishing gradients and inefficiency in long-sequence processing. Transformers [42, 43, 44] excel at capturing long-range temporal dependencies and demonstrate potential, though their efficiency in real-time applications remains unresolved. Existing methods often modify base network architectures to improve performance. For instance, the Temporal Segment Network (TSN) [45] divides videos into segments for feature extraction and classification, effectively capturing temporal dynamics. This approach has been applied to action classification in the Assembly101 dataset, demonstrating advantages in modeling temporal information.
@@ 2.1 Egocentric Gesture Recognition
2.2 Event-based Methods for Human Perception
Event cameras, inspired by biological vision systems, asynchronously record pixel intensity changes, making them highly suitable for dynamic scenes with fast motion [10, 12, 46]. Recent works have explored the use of event cameras for human perception tasks such as human pose estimation [47, 16, 48, 49], action classification [50, 51], eye tracking [52, 53, 54, 55], sign language [56, 57] and head pose estimation [58], demonstrating their advantages in temporal resolution and robustness to motion blur. All existing event-based studies on hand motion analysis are conducted from a third-person perspective. DVS128 [1] proposed the first event-based hardware system for gesture recognition, achieving high accuracy and low latency under low-power conditions for gesture recognition. However, this system faces challenges such as low resolution and a lack of hand pose diversity. EventHands [13] achieves 1kHz 3D hand pose estimation using event cameras, which leverages Locally Normalized Event Surfaces (LNES) to aggregate asynchronous event data into a compact 2D representation, emphasizing temporal locality and enabling efficient real-time processing with high accuracy. This method has been widely applied in the field of human body perception [59, 16], achieving excellent results. EvHandPose [60] employs a novel hand flow representation and a weakly-supervised framework to address motion ambiguities in event streams, though its focus lies in reconstruction rather than gesture recognition. Ev2Hands [61] achieved the first 3D tracking of two fast-moving and interacting hands from a single monocular event camera. EHoA [62] pioneered hand-object action recognition via event vision. EvRGBHand [63] introduced complementary use of event streams and RGB frames for hand mesh reconstruction. E2(GO)MOTION [64] synthesized the N-EPIC-Kitchens dataset, the first egocentric action recognition dataset based on event data. Their work highlights the potential of event data for action recognition, particularly excelling in scenarios with rapid motion and low lighting. The recent work Helios [59] proposes an event-based gesture recognition system optimized for smart eyewear. The study introduces a synthetic event dataset with seven gesture classes to train the recognition model, showcasing the potential for low-power, real-time implementation in wearable devices. However, the system exhibits notable accuracy degradation in real-world testing compared to synthetic data performance, and the restricted gesture taxonomy limits its generalizability. Meanwhile, the authors do not address the open-source accessibility of the dataset. However, reliance on synthetic data introduces challenges due to the real-synthetic domain gap.
The unique asynchronous and sparse nature of event data creates inherent incompatibilities with conventional RGB-based processing paradigms, which are designed for synchronous frame-based inputs. While event streams offer microsecond temporal resolution and sparse asynchronous encoding of luminance changes, their ultra-high temporal granularity introduces unique computational challenges. Existing event-based approaches predominantly focus on instantaneous feature extraction tasks (e.g., 3D keypoint estimation) rather than continuous temporal understanding. This limitation stems from two inherent properties: (1) The asynchronous discrete nature of event representation complicates the modeling of coherent temporal dynamics, and (2) The fragmented temporal cues inherent in sparse event streams hinder process-level semantic interpretation. Consequently, current methodologies struggle to reconcile the conflict between leveraging fine-grained motion dynamics captured by event sensing and establishing holistic temporal reasoning frameworks. This modality discrepancy underscores the critical need for dedicated architectures addressing temporal continuity in event-based gesture recognition.
A comprehensive comparison of existing gesture recognition datasets across modalities is presented in Table 1. The table systematically evaluates 18 datasets through eight critical dimensions: sensing modality, data volume, action classes, two-hand interaction support, spatial resolution, publication year, real-world applicability, and first-person perspective. Notably, EgoEvGesture (our proposed dataset) establishes four key advantages: (1) First and only event-based dataset for egocentric gestures – Existing event datasets focus on third-person actions, while ours captures head-motion-contaminated events unique to first-person view. (2) the largest event-based sample volume (5,419 samples), (3) the richest gesture vocabulary among event-based datasets (38 classes), and (4) native support for two-hand interaction in egocentric scenarios - features previously unavailable in existing event-stream benchmarks.
Table 1: A comprehensive comparison of relative datasets including details on their modalities, volumes, action classes, two-hand interaction presence, resolutions, years of publication, real-world applicability, and first-person perspective availability.
@@ 2.2 Event-based Methods for Human Perception
1st-P
Ego-Gesture [7] RGB 3M 83 640p 18
FPHA [24] RGB 0.1M 45 1080p 18
EPIC-KITCHENS [23] RGB 11.5M 149 1080p 18
H20 [25] RGB 0.57M 36 720p 21
Assembly101 [26] RGB 110M 1380 1080p 21
ARCTIC [27] RGB 2.1M 11 2.8k 24
HOI4D [28] RGB 2.4M 54 8k 24
DVS128 [1] Event 1342S 10 128² 17
N-EPIC [64] Event NA 8 225² 22
EHoA [62] E+R 2144S 8 346p 24
EvRealHands [60] E+R 79m NA 720p 23
Ev2Hands-S [61] Event * NA 512² 23
Ev2Hands-R [61] E+R 20m NA 346² 23
EventHands [13] Event 100h NA 240p 21
EventEgo3D [16] Event 127m NA 320² 24 Both
Helios [59] Event 600S 7 320² 24
EVREALHANDS [63] E+R 74m NA 346p 24
EgoEvGesture (Ours) Event 5419S 38 720p 25
✓ = Yes  ✗ = No Abbreviations: “E+R” denotes Event+RGB, “2H” denotes Two-Hand Interaction, “Cls” denotes Classes, “NA” denotes Not Available
Units: “M” denotes Million, “S” denotes Samples, “m” denotes minutes, “h” denotes Hours, denotes Number of events (3.12 $\times$ 108)
3 EgoEvGesture Dataset
3.1 Head-Mounted Capture System
Our egocentric gesture acquisition platform (illustrated in Figure 1(a)) employs a custom-designed head-mounted system (HMS) comprising a lightweight bicycle helmet modified with 3D-printed mounts to rigidly affix two Prophesee EVK4 event cameras and an Intel RealSense D435 depth camera. Temporal synchronization between event streams and RGB-D frames is achieved through hardware triggering, where the Intel RealSense D435 emits a synchronization signal to align data acquisition across all sensors. The event cameras utilize 4 mm focal length lenses (MV-LD-4-4M-G) providing a 94° horizontal field of view (FOV) with a native resolution of $1280\times 720$ pixels. Each pixel has a size of $4.86\mu m\times 4.86\mu m$ with temporal resolution up to 1 MHz, capturing brightness changes through asynchronous events. For normal light conditions, we employed the default EVK4 threshold configuration (0% bias_diff_off and 0% bias_diff_on). In low-light scenarios, sensitivity was enhanced by configuring 20% bias_diff_off and 20% bias_diff_on parameters to increase event generation rate. The Intel RealSense D435 RGB-D camera operates at 30 fps with synchronized $1280\times 720$ resolution for depth streams, which may be used for further research. The system connects via USB 3.0 to a host computer for power delivery and real-time data streaming, enabling uninterrupted recording sessions. The compact design and flexible head-mounted system (HMS) allow users to freely move their heads and perform hand movements while ensuring precise temporal alignment.
(a) Proposed head-mounted capture system.
(b) Recording environment.
Figure 1: (a) Proposed head-mounted capture system. It integrates two Prophesee EVK4 event cameras (left/right) and an Intel RealSense D435 RGB-D sensor (center). (b) Recording environment. Each participant is equipped with the HMS and performed a series of 38 distinct gestures under normal light and low light conditions.
3.2 Data Collection Protocol
The dataset consists of data from 10 participants, encompassing a diverse range of body types and genders. Each participant is equipped with the HMS and performed a series of 38 distinct gestures (as illustrated in Figure 2) under both normal and low-light conditions(as illustrated in Figure 1(b)). Each gesture is executed at least four times by each participant, ensuring a robust and varied dataset. The gestures are categorized into three main groups:
Figure 2: The 38 classes of hand gesture designed in our proposed EgoEvGesture dataset. The figure shows events represented by LNES, with arrows indicating the direction of hand (thick arrows) or finger (thin arrows) movements. The freestyle motions vary among participants, with three examples shown in the figure.
  • Single-Hand Gestures (Right Hand): 11 classes including Wave to the right, Wave to the left, Right thumb points to the right, Right index finger points to the left, Tap, Spread, Shrink, Swipe, Draw a circle, Draw an X, NO (Shake finger).
  • Two-Hand Gestures (No Occlusion): 21 classes including Wave down, Wave up, Wave normally (arms close and move away), Raise arms, Push forward, OK, Flip palms, Grab, Make a heart, Camera gesture, Thumbs up, Show number 1-9, Make a fist.
  • Two-Hand Gestures (Mutual Occlusion): 6 classes emphasizing challenging hand-hand interactions including Interlock fingers, Clap hands, Cross index fingers, Rotate arms, Salute with a fist, Freestyle motion.
Each participant performed all gestures under controlled normal illumination and low light conditions, with at least four repetitions per gesture-condition combination. Figure 3 contrasts the outputs of RGB and event cameras across lighting regimes, demonstrating the robustness of the latter to illumination variations.
@@ 3.2 Data Collection Protocol
Figure 3: A comparison of what an event camera (top) and an RGB camera (bottom) captured under normal light (left) and low light (right) conditions. Event cameras are almost unaffected by lighting conditions, while RGB images are barely distinguishable in low light conditions.
3.3 Dataset Composition and Statistics
[…]

### v2

@@
Abstract
Egocentric gesture recognition is a pivotal technology for enhancing natural human-computer interaction, yet traditional RGB-based solutions suffer from motion blur and illumination variations in dynamic scenarios. While event cameras show distinct advantages in handling high dynamic range with ultra-low power consumption, existing RGB-based architectures face inherent limitations in processing asynchronous event streams due to their synchronous frame-based nature. Moreover, from an egocentric perspective, event cameras record data that includes events generated by both head movements and hand gestures, thereby increasing the complexity of gesture recognition. To address this, we propose a novel network architecture specifically designed for event data processing, incorporating (1) a lightweight CNN with asymmetric depthwise convolutions to reduce parameters while preserving spatiotemporal features, (2) a plug-and-play state-space model as context block that decouples head movement noise from gesture dynamics, and (3) a parameter-free Bins-Temporal Shift Module (BSTM) that shifts features along bins and temporal dimensions to fuse sparse events efficiently. We further establish the EgoEvGesture dataset, the first large-scale dataset for egocentric gesture recognition using event cameras. Experimental results demonstrate that our method achieves 62.7% accuracy tested on unseen subjects with only 7M parameters, 3.1% higher than state-of-the-art approaches. Notable misclassifications in freestyle motions stem from high inter-personal variability and unseen test patterns differing from training data. Moreover, our approach achieved a remarkable accuracy of 97.0% on the DVS128 Gesture, demonstrating the effectiveness and generalization capability of our method on public datasets. The dataset and models are made available at https://github.com/3190105222/EgoEv_Gesture.
I Introduction
(a) Proposed head-mounted capture system.
(b) Recording environment (HDR).
(c) Representation of 38 gesture classes.
Figure 1: (a) Proposed head-mounted capture system (HMCS). It integrates two Prophesee EVK4 event cameras (left/right) and an Intel RealSense D435 RGB-D sensor (center). (b) Recording environment. Each participant is equipped with the HMCS and performs a series of 38 distinct gestures under normal light and low light conditions. (c) The 38 classes of hand gesture designed in our established EgoEvGesture dataset. The figure shows events visualization, with arrows indicating the direction of hand (thick arrows) or finger (thin arrows) movements. The freestyle motions vary among participants, with three examples shown in the figure.
@@ I Introduction
To address these challenges, we first propose a novel approach with three progressive processing stages, as illustrated in Figure 5. First, we convert raw events into Locally Normalized Event Surfaces (LNES) [12] that focus on the most recent temporal signature, suppressing historical ego-motion artifacts while maintaining real-time performance. Second, we employ Blaze [13] and Vision Mamba (VMamba) [14] blocks to extract spatiotemporal features. The Blaze encoder-decoder, adapted from [13] with proven efficiency in human-centric perception [15], processes each temporal bin efficiently. The VMamba block  [14] addresses sparse event processing and long-term temporal modeling through discrete state space equations. Finally, we propose the Bins-Temporal Shift Module (BTSM), a zero-parameter mechanism that shifts features bidirectionally along two orthogonal axes: 1) Bins axis (fine temporal resolution: 33 ms/bin) for inter-bin context fusion, 2) Raw temporal axis (coarse resolution: 200 ms) for intra-bin motion alignment. This dual-scale shifting resolves temporal fragmentation while maintaining motion invariance.
Due to the lack of event datasets in the egocentric setting, we developed EgoEvGesture, the first large-scale dataset for egocentric gesture recognition using event cameras. Captured under both normal and low-light conditions, it includes data from 10 subjects performing a variety of gestures, offering rich temporal and spatial information for training and evaluating models in real-world scenarios. Our proposed method achieves a classification accuracy of 62.7% on this dataset, with only 7M parameters, outperforming existing (closely related) approaches by 3.0%. To verify the effectiveness of our approach, we evaluated it on the DVS128 Gesture dataset [16], achieving a competitive accuracy of 97.0%. This result highlights the robustness of our architecture in handling sparse and temporally challenging event-based data. The significance of each proposed module is evaluated and confirmed in an ablation study.
In summary, this paper defines a new problem, i.e., gesture recognition based on an egocentric event camera, and makes the following technical contributions:
@@
II Related Work
Next, we review related works on egocentric gesture recognition and event-based approaches to human perception. As shown in Table I, our EgoEvGesture represents the first dataset for egocentric gesture recognition that uses real-world event streams, offering several advantages over existing RGB- or depth-based methods, as highlighted in the following sections.
II-A Egocentric Gesture Recognition
Egocentric gesture recognition using RGB and depth sensors has been extensively explored [17, 18, 5, 19]. The EgoGesture dataset [6], one of the largest benchmarks, contains 83 gesture classes designed for wearable device interactions, which inspired our motion design. The EPIC-KITCHENS dataset [20] focuses on action classification in kitchen environments, but prioritizes general activities over specific gestures. The FPHA dataset [21] pioneered 3D hand pose annotations in egocentric scenarios but suffers from magnetic sensor interference and lacks two-hand pose data. Several datasets have advanced egocentric hand-object interaction research, yet each has limitations for dynamic gesture recognition. The H2O  [22] and Assembly101 [23] offer multi-view videos but lacks gesture diversity, while ARCTIC [24] provides dexterous bimanual manipulation data but lacks dynamic gesture annotations. HOI4D [25] enriches single-hand interaction with 3D meshes and segmentation but omits two-hand scenarios.
@@ II-A Egocentric Gesture Recognition
II-B Event-based Methods for Human Perception
TABLE I: Ultra-compact dataset comparison with key attributes
@@ II-B Event-based Methods for Human Perception
1P
Ego-Gesture [6] R 3M 83 $\checkmark$ 640 18 $\checkmark$ $\checkmark$
FPHA [21] R 0.1M 45 1080 18 $\checkmark$ $\checkmark$
EPIC-K [20] R 11M 149 $\checkmark$ 1080 18 $\checkmark$ $\checkmark$
H20 [22] R 0.6M 36 $\checkmark$ 720 21 $\checkmark$ $\checkmark$
Asm101 [23] R 110M 1k 1080 21 $\checkmark$ $\checkmark$
ARCTIC [24] R 2M 11 $\checkmark$ 2.8k 24 $\checkmark$ $\checkmark$
HOI4D [25] R 2M 54 8k 24 $\checkmark$ $\checkmark$
DVS128 [16] E 1kS 10 128 17 $\checkmark$
N-EPIC [26] E NA 8 $\checkmark$ 225 22 $\checkmark$
EHoA [27] E+R 2kS 8 346 24 $\checkmark$
EvRealH [28] E+R 79m NA 720 23 $\checkmark$
Ev2H-S [29] E * NA $\checkmark$ 512 23
Ev2H-R [29] E+R 20m NA $\checkmark$ 346 23 $\checkmark$
EvHands [12] E 100h NA 240 21
EvEgo3D [15] E 127m NA 320 24 B $\checkmark$
Helios [30] E 600S 7 320 24 $\checkmark$ $\checkmark$
EvRealH2 [31] E+R 74m NA 346 24 $\checkmark$
Ours E 5kS 38 $\checkmark$ 720 25 $\checkmark$ $\checkmark$
$\checkmark$=Yes, –=No, B=Both.
Attributes: M: Modality (R=RGB, E=Event), V: Volume, C: Classes, 2H: Two-Hand, R: Resolution (pixels), Y: Year, W: Real-World, 1P: 1st-Person
Units: M=106, k=103, S=Samples, m=minutes, h=hours, *=3.1$\times$108 events
Event cameras, inspired by biological vision systems, asynchronously record pixel intensity changes, making them highly suitable for dynamic scenes with fast motion [9]. Recent works have explored the use of event cameras for human perception tasks such as human pose estimation [15], action classification [16], eye tracking [32], sign language recognition [33], and head pose estimation [34], demonstrating their advantages in temporal resolution and robustness to motion blur. Most existing event-based studies on hand motion analysis are conducted from a third-person perspective. DVS128 [16] proposed the first event-based hardware system for gesture recognition. However, this system faces challenges such as low resolution and a lack of hand pose diversity. EventHands [12] proposes Locally Normalized Event Surfaces (LNES) to aggregate asynchronous event data into a compact 2D representation and achieves 1kHz 3D hand pose estimation. EvHandPose [28] employs a novel hand flow representation and a weakly-supervised framework to address motion ambiguities in event streams, though its focus lies in reconstruction rather than gesture recognition. Ev2Hands [29] achieved the first 3D tracking of two fast-moving and interacting hands from a single monocular event camera. EHoA [27] pioneered hand-object action recognition via event vision. EvRGBHand [31] introduced complementary use of event streams and RGB frames for hand mesh reconstruction. E2(GO)MOTION [26] synthesizes the N-EPIC-Kitchens dataset, the first egocentric action recognition dataset based on event data. The recent work Helios [30] proposes an event-based gesture recognition system optimized for smart eyewear, trained on a synthetic 7-class dataset to demonstrate low-power real-time operation. However, it shows a significant real-world accuracy drop, limited gesture diversity, and lacks dataset openness, while synthetic training introduces domain gap challenges.
The unique asynchronous and sparse nature of event data conflicts with conventional synchronous frame-based RGB processing. Despite microsecond temporal resolution, this ultra-high temporal granularity introduces unique computational challenges. Existing event-based approaches predominantly focus on instantaneous feature extraction tasks rather than continuous temporal understanding. This limitation stems from two inherent properties: (1) The asynchronous discrete nature of event representation complicates the modeling of coherent temporal dynamics, and (2) The fragmented temporal cues inherent in sparse event streams hinder process-level semantic interpretation. Consequently, current methodologies struggle to reconcile the conflict between leveraging fine-grained motion dynamics captured by event sensing and establishing holistic temporal reasoning frameworks. This modality discrepancy underscores the critical need for dedicated architectures addressing temporal continuity in event-based gesture recognition.
A comprehensive comparison of existing gesture recognition datasets across modalities is presented in Table I. The table systematically evaluates 18 datasets through eight critical dimensions: sensing modality, data volume, action classes, two-hand interaction support, spatial resolution, publication year, real-world applicability, and first-person perspective. Notably, EgoEvGesture (our proposed dataset) establishes four key advantages: (1) First and only event-based dataset for egocentric gestures – Existing event datasets focus on third-person actions, while ours captures head-motion-contaminated events unique to first-person view. (2) the largest event-based sample volume (5,419 samples), (3) the richest gesture vocabulary among event-based datasets (38 classes), and (4) native support for two-hand interaction in egocentric scenarios - features previously unavailable in existing event-stream benchmarks.
@@
Figure 2: A comparison of what an event camera (top) and an RGB camera (bottom) captured under normal light (left) and low light (right) conditions. Event cameras are almost unaffected by lighting conditions, while RGB images are barely distinguishable in low-light conditions.
III EgoEvGesture Dataset
III-A Head-Mounted Capture System (HMCS)